# Confidence-Aware Multi-Field Model Calibration

### v1

@@ Abstract.
Accurately predicting the probabilities of user feedback, such as clicks and conversions, is critical for advertisement ranking and bidding. However, there often exist unwanted mismatches between predicted probabilities and true likelihoods due to the shift of data distributions and intrinsic model biases. Calibration aims to address this issue by post-processing model predictions, and field-aware calibration can adjust model output on different feature field values to satisfy fine-grained advertising demands. Unfortunately, the observed samples corresponding to certain field values can be seriously limited to make confident calibrations, which may yield bias amplification and online disturbance. In this paper, we propose a confidence-aware multi-field calibration method, which adaptively adjusts the calibration intensity based on the confidence levels derived from sample statistics. It also utilizes multiple feature fields for joint model calibration with awareness of their importance to mitigate the data sparsity effect of a single field. Extensive offline and online experiments show the superiority of our method in boosting advertising performance and reducing prediction deviations.
@@ 1. Introduction
Pioneering studies like NeuCalib (Pan et al., 2020) and AdaCalib (Wei et al., 2022) have focused on field-aware calibration by learning additional calibration neural networks to adjust original model responses. However, the samples corresponding to certain feature values can often be too sparse to learn an unbiased and confident calibrating network. As shown in Fig. 1, the calculated CTR is 0.1 if an ad slot accumulates 50 clicks out of 500 impressions, while the 95% confidence interval of CTR is [0.076, 0.129] (calculated by the Wilson interval formula (Wilson, 1927)). Such a wide range indicates strong uncertainty in field-aware calibration. Furthermore, according to our industrial practice, the volumes of click and impression data w.r.t. most field values are smaller than the example above, which poses a severe challenge to the confidence of calibration models. In addition, different fields may have diverse probability score distributions (Fig. 1 left), and ad platforms may select multiple fields as the business targets. Most existing methods mainly focus on single field calibration, and it is impractical to directly merge multiple fields due to the ultra sparsity of samples satisfying multiple field value conditions.
In this paper, we propose a confidence-aware multi-field calibration method named ConfCalib, which can perform fine-grained calibration on multiple feature fields meanwhile keeping awareness of the confidence degree of field data observation to adjust the calibration intensity adaptively. Compared with methods based on neural networks, ConfCalib, as a non-parametric method, performs more robustly in scenarios with severe data sparsity and has low costs of computation and deployment. Concretely, we regard user feedback on each field value as a binomial distribution and use Wilson intervals to compute confidence-aware deviation degrees of model predictions, where larger deviations indicate greater discrepancies between observed and predicted probabilities. We devise a mapping function to derive a smaller deviation and transform the predicted scores to achieve this new deviation. To further eliminate the influence of data sparsity, we develop a simple yet effective multi-field fusion method to synthesize the calibration results on different fields to enhance calibration performance and better accommodate advertising demands. Extensive offline experiments in different scenarios and the online A/B test on our advertising platform demonstrate the effectiveness of our approach.
@@ 2. Related Work
Model calibration is a classic problem in machine learning, which is proposed to eliminate the discrepancy between the predicted probabilities of classification models and the true likelihood (Wald et al., 2021; Pleiss et al., 2017; Nixon et al., 2019). It has been widely explored due to the requirement of accurate probability prediction in various applications, such as ad auctions (He et al., 2014; McMahan et al., 2013), weather forecasting (Gneiting and Raftery, 2005; DeGroot and Fienberg, 1983), and personalized medicine (Jiang et al., 2012).
Calibration methods calibrate trained prediction models during training or inference phase. Methods integrated into training usually conduct calibration through well-designed loss functions, including AvUC (Krishnan and Tickoo, 2020), MMCE (Kumar et al., 2018), Focal loss (Mukhoti et al., 2020; Lin et al., 2017) and SB-ECE (Karandikar et al., 2021), etc. Besides, label smoothing, which can be interpreted as a modified primary loss, also serves to model calibration (Müller et al., 2019; Pereyra et al., 2017; Liu et al., 2022). For model calibration in recommendation systems, post-hoc calibration methods are more prevalent. These methods transform the output of trained prediction models in the post-processing phase. According to the characteristics of different mapping functions, existing calibration methods can be divided into three categories: scaling-based, binning-based, and hybrid.
@@
Hybrid methods (Kumar et al., 2019; Zhang et al., 2020; Deng et al., 2021; Huang et al., 2022; Pan et al., 2020; Wei et al., 2022) combine scaling and binning to fit various mapping functions for different sample subsets, which allows a more precise calibration than single scaling or binning methods. Scaling-binning (Kumar et al., 2019) first fits a mapping function for the sorted samples and then employs histogram binning calibration on the mapped scores. Smoothed Isotonic Regression (SIR) (Deng et al., 2021) uses isotonic regression on bins and fits a monotonic piece-wise linear calibration function. Multiple Boosting Calibration Trees (MBCT) (Huang et al., 2022) optimizes the binning scheme through a tree-like division along feature fields and adopts linear scaling on each tree node individually. Recently, some works have focused on field-wise calibration, which aims to reduce the calibration error in a specific feature field. Neural Calibration (Pan et al., 2020) learns a neural isotonic line-plot scaling function and an auxiliary network directly taking features as input to capture feature field information. AdaCalib (Wei et al., 2022) extends the single global mapping function of Neural Calibration to different ones for each value of a certain feature field and dynamically adjusts the number of bins by neural networks. These neural methods may suffer from data sparsity in specific application scenarios because networks are difficult to train. Additionally, existing field-wise calibration methods simply consider one field and ignore potential relations among fields, meaning there is still a lack of research on multi-field calibration methods.
@@ 3. Methodology
The overall framework of our multi-field confidence-aware calibration method is shown in Fig. 2. It first performs confidence-aware calibration on multiple target feature fields independently and then fuses the scores calibrated on different feature fields into a unified one. The details of our approach are introduced as follows.
@@ 3.2. Confidence-aware Calibration
Next, we introduce our confidence-aware calibration method. In practical scenarios, field-level calibration faces severe data sparsity issues due to the limited time window for data collection and the intrinsic sparsity of certain user feedback, such as conversions. Consequently, calibrations on fields with numerous feature values are neither accurate nor confident. The main problem caused by data sparsity is that the observed probability is not credible enough to reflect the true data distribution. In this case, complete trust in limited observations will likely result in overfitting. Therefore, it is necessary to preserve some trust in model predictions while referencing observations under data sparsity.
Actually, many types of user feedback, like clicks and conversions, can be naturally modeled as binomial distributions. Given the observed positive sample ratio $p$ and the number of total samples $n$, we can then use a confidence interval formula to compute the confidence interval ($p_{-}$, $p_{+}$) of the real feedback probability under a certain confidence level. Conversely, given a certain predicted feedback probability, we can obtain the corresponding confidence level by considering the predicted value as the bound of the confidence interval. In other words, there is a mapping between the predicted probability and the confidence level, which is the core of our confidence-aware calibration method.
The most commonly used confidence interval formula is the normal interval. However, due to the limitation of the central limit theorem, it is unreliable with a small sample size or an extreme probability close to 0 or 1, which is quite common for user feedback prediction in recommender systems. Therefore, we adopt the Wilson confidence interval (Wilson, 1927), which can be safely employed under the special circumstances above. The formula is
@@
In order to perform field-level calibration, our approach operates at the level of sample subsets, which can be obtained by splitting samples based on different values of the target feature field. Besides, field-wise binning [^1] Binning is optional in our method since it introduces heavier data sparsity., i.e., further dividing each subset into several bins by order of their predicted scores, is optional. For each sample subset, we need the statistics, including the number of samples $n$, the ratio of positive samples $p$, and the average predicted score $\hat{p}$. By substituting these variables into Eq. (1), we can solve the value of the deviation score $z\in[0,+\infty)$ using numerical algorithms, such as the bisection method. A larger $z$ indicates greater discrepancies between model predictions and current observations. To calibrate the predictions, we propose to use a non-linear transformation function $g(z)$ to reduce the deviation. Intuitively, this function should satisfy the following requirements: (1) passing through the origin: $g(0)=0$; (2) monotonic: $\forall z_{1}\leq z_{2},g(z_{1})\leq g(z_{2})$; (3) bounded: $\exists C\in\mathbb{R}^{+},g(+\infty)\leq C$. Thus, we devise a variant of the sigmoid function as follows:
@@
where $\lambda$ is a hyperparameter used to provide an upper bound and control the calibration intensity manually for large z-scores, i.e., deviations. A large calibration intensity means more belief in observations and less belief in original predictions.
Obviously, we have $g(+\infty)=\lambda$ so that large deviations are calibrated to a relatively small one (e.g., the deviation corresponding to 95% confidence level if $\lambda=1.96$), and minor deviations are reduced almost linearly. By replacing the original $z$ in Eq. (1) with the transformed value $g(z)$, we can obtain the calibrated average probability score $\hat{p}^{\prime}$. In order to ensure that $\hat{p}^{\prime}$ is located between the observed probability $p$ and the predicted probability $\hat{p}$, we use the positive branch if $\hat{p}$ is greater than $p$. Otherwise, we use the negative one. In this way, given a test sample located in the same subset, we compute a field-wise calibrated score $p_{\text{calib}}$ by scaling its original predicted probability $p_{\text{pred}}$, i.e.,
@@
Overall, our confidence-aware method adjusts model-predicted probabilities by resetting the confidence level, which considers the observation and prediction together to avoid violent value changes and overfitting limited observations. Consequently, it keeps robust when handling severe sample sparsity, such as circumstances with few observed samples and zero positive samples.
@@ 3.3. Multi-field Joint Calibration
Previous field-level calibration methods only calibrate and evaluate errors in a certain field, ignoring other fields. Actually, the field-level calibration error in different fields may vary greatly even for the same dataset, indicating a natural discrepancy of prediction biases in different fields. In addition, the biases in different fields, as reflections of the global bias from different perspectives, have a certain degree of commonality. Therefore, the data sparsity problem faced by a certain field is expected to be alleviated through calibrations on other fields. Therefore, simultaneously capturing biases of multiple fields can be significant to eliminate overall prediction bias and improve calibration performance.
A naive way to extend calibration from a single field to multiple fields is to split sample subsets based on the combination of feature values of all the fields. Nevertheless, it is infeasible for the severer data sparsity since fewer samples satisfy multiple field value conditions. This means prior field-wise calibration methods for a single field cannot be directly extended to multi-field calibration.
On the contrary, our method demonstrates better extensibility among fields. Note that the confidence-aware calibration of a certain field in Section 3.2 is actually a scaling operation for a specific sample and thus can be easily generalized to a multi-field joint calibration method. Assume that there are $K$ target fields to calibrate. Given a sample, we will get $K$ scaling multipliers $m_{1},m_{2},\dots,m_{K}$, i.e. ${\hat{p}^{\prime}}/{\hat{p}}$ of each field. Then, a natural manner to adaptively fuse these multipliers is the weighted geometric mean,
@@
where $w_{1},w_{2},\dots,w_{K}$ is weights satisfying $\sum_{i=1}^{K}{w_{i}}=1$. Considering the relatively small computation cost, the fusion weights can be precisely obtained by grid search. This method allows us to conduct joint calibration based on the marginal distribution of each field independently instead of dealing with the joint distribution, thus avoiding further sparsity of samples.
@@ 3.4. Online Deployment
Fig. 4 shows the online deployment diagram of ConfCalib in our advertising platform. On the offline side, the prediction model’s training and ConfCalib’s fitting are paralleled and completely independent. At fixed intervals, the calibration module pulls the latest online samples from the log database and refits calibration curves. This interval is usually set relatively shorter (e.g., 30 minutes) than the prediction model update period to better capture the dynamic shift of online data. On the online side, prediction and calibration are conducted in sequence. Firstly, the prediction model makes predictions for candidate objects, then the calibration model calibrates the predicted probabilities and outputs the calibrated probabilities for subsequent use in the platform.
@@ 4.1.1. Settings
For all the datasets and methods, the prediction model is a fully connected network with four layers. Hyperparameters of comparative methods, like the number of bins, are set according to the chosen value in the corresponding papers if provided. Otherwise, we tune the hyperparameters and select the optimal values. Considering the strong randomness of neural methods, we conduct five random experiments on each of them and report means and standard deviations of the results. For our method we set the number of bins as {50, 1, 10} and $\lambda$ in the z-score mapping function as {1.0, 0.2, 1.6}, corresponding to the following three datasets.
Table 1. Detailed statistics of the three datasets.

### v2

@@ Abstract.
Accurately predicting the probabilities of user feedback, such as clicks and conversions, is critical for advertisement ranking and bidding. However, there often exist unwanted mismatches between predicted probabilities and true likelihoods due to the rapid shift of data distributions and intrinsic model biases. Calibration aims to address this issue by post-processing model predictions, and field-aware calibration can adjust model output on different feature field values to satisfy fine-grained advertising demands. Unfortunately, the observed samples corresponding to certain field values can be seriously limited to make confident calibrations, which may yield bias amplification and online disturbance. In this paper, we propose a confidence-aware multi-field calibration method, which adaptively adjusts the calibration intensity based on confidence levels derived from sample statistics. It also utilizes multiple fields for joint model calibration according to their importance to mitigate the impact of data sparsity on a single field. Extensive offline and online experiments show the superiority of our method in boosting advertising performance and reducing prediction deviations.
@@ 1. Introduction
Pioneering studies like NeuCalib (Pan et al., 2020) and AdaCalib (Wei et al., 2022) have focused on field-aware calibration by learning additional calibration neural networks to adjust original model predictions. However, samples corresponding to certain feature values are often too sparse to learn an unbiased and confident calibrating network. As shown in Fig. 1, the calculated CTR is 0.1 if an ad slot accumulates 50 clicks out of 500 impressions, while the 95% confidence interval of CTR is [0.076, 0.129] (calculated by the Wilson interval formula (Wilson, 1927)). Such a wide range indicates strong uncertainty in field-aware calibration. Furthermore, according to our industrial practice, the volumes of click and impression data w.r.t. most field values are smaller than the example above, which poses a severe challenge to the confidence of calibration models. In addition, different fields may have diverse probability score distributions (Fig. 1 left), and ad platforms may select multiple fields as the business targets. Most existing methods mainly focus on single field calibration, and it is impractical to directly merge multiple fields due to the ultra sparsity of samples satisfying multiple field value conditions.
In this paper, we propose a confidence-aware multi-field calibration method named ConfCalib, which can perform fine-grained calibration on multiple feature fields and meanwhile keep robustness under data sparsity via adaptive adjustment of the confidence level. Concretely, we assume that user feedback follows a binomial distribution and use Wilson intervals to compute confidence-aware calibrated scores with a dynamic adjustment of calibration intensity. Moreover, we propose a simple yet effective multi-field fusion method to synthesize the calibration results on different fields, in order to further eliminate the influence of data sparsity. Compared with methods based on neural networks, our ConfCalib performs more robustly in scenarios with severe data sparsity. Extensive offline experiments in different datasets and online A/B test on our advertising platform demonstrate the effectiveness of our method.
@@ 2. Related Work
Model calibration is a classic problem in machine learning, which is proposed to eliminate the discrepancy between the predicted probabilities of classification models and the true likelihood (Wald et al., 2021; Pleiss et al., 2017; Nixon et al., 2019). It has been widely explored to promote accurate probability prediction in various applications, such as ad auctions (He et al., 2014; McMahan et al., 2013), weather forecasting (Gneiting and Raftery, 2005; DeGroot and Fienberg, 1983), and personalized medicine (Jiang et al., 2012).
Calibration methods calibrate trained prediction models during training or inference phase. Methods integrated into training usually conduct calibration through well-designed loss functions, including AvUC (Krishnan and Tickoo, 2020), MMCE (Kumar et al., 2018), Focal loss (Mukhoti et al., 2020; Lin et al., 2017) and SB-ECE (Karandikar et al., 2021), etc. Besides, label smoothing, which can be interpreted as a modified primary loss, also serves to model calibration (Müller et al., 2019; Pereyra et al., 2017; Liu et al., 2022). This paper focus on post-hoc methods that calibrate model predictions by post-processing. According to the characteristics of different mapping functions, existing calibration methods can be divided into three categories: scaling-based, binning-based, and hybrid methods.
@@
Hybrid methods (Kumar et al., 2019; Zhang et al., 2020; Deng et al., 2021; Huang et al., 2022; Pan et al., 2020; Wei et al., 2022) combine scaling and binning methods, which are superior in calibration performance and prevalence in industry. Scaling-binning (Kumar et al., 2019) first fits a mapping function for the sorted samples and then employs histogram binning calibration on the mapped scores. Smoothed Isotonic Regression (SIR) (Deng et al., 2021) uses isotonic regression on bins and fits a monotonic piece-wise linear calibration function. Recently, some works have focused on field-wise calibration, which aims to reduce the calibration error in a specific feature field. Neural Calibration (Pan et al., 2020) learns a neural isotonic line-plot scaling function and an auxiliary network directly taking features as input to capture feature field information. AdaCalib (Wei et al., 2022) extends the single global mapping function of Neural Calibration to different ones for each value of a certain feature field and dynamically adjusts the number of bins by neural networks. These neural methods may suffer from data sparsity in specific application scenarios where network training is hard to converge. Additionally, existing field-wise calibration methods simply consider one field and ignore potential relations among fields, meaning there is still a lack of research on multi-field calibration methods.
@@ 3. Methodology
Figure 2. The overall framework of our ConfCalib method.
The overall framework of our multi-field confidence-aware calibration method is shown in Fig. 2. It first performs confidence-aware calibration on multiple target feature fields independently and then fuses the scores calibrated on different feature fields into a unified one. The details of our method are introduced as follows.
@@ 3.2. Confidence-aware Calibration
Next, we introduce our confidence-aware calibration method. In practical scenarios, field-level calibration faces severe data sparsity issues due to the limited time window for data collection and intrinsic sparsity of certain user feedback, such as conversions. Consequently, calibrations on feature fields with numerous values are neither accurate nor confident. The main problem caused by data sparsity is that the observed probability is not credible enough to reflect the true data distribution. In this case, complete trust in limited observations will result in overfitting and harm generalization. In summary, it is necessary to preserve some trust in model predictions while referencing observations under data sparsity.
Most user feedback, like clicks and conversions, can be naturally modeled using binomial distributions. Given the observed positive sample ratio $p$ and the number of total samples $n$, we can use a confidence interval formula to compute the confidence interval (CI) ($p_{-}$, $p_{+}$) of the real feedback probability under a certain confidence level. Conversely, given a certain predicted feedback probability, we can obtain the corresponding confidence level by regarding the predicted value as the bound of CI. In other words, there is a mapping between the predicted probability and the confidence level, which is the core of our confidence-aware calibration method.
The most commonly used CI formula is the normal interval. However, due to the limitation of the central limit theorem, it is unreliable with a small sample size or an extreme probability close to 0 or 1, which is quite common for user feedback prediction in recommender systems. Therefore, we adopt the Wilson confidence interval (Wilson, 1927), which can be safely employed under the special circumstances above. The formula is
@@
In order to perform field-level calibration, our method operates at the level of sample subsets, which can be obtained by splitting samples based on different values of the target feature field. Besides, field-wise binning [^1] Binning is optional in our method since it introduces heavier data sparsity., i.e., further dividing each subset into several bins by order of their predicted scores, is optional. For each sample subset, we need the statistics, including the number of samples $n$, the ratio of positive samples $p$, and the average predicted score $\hat{p}$. Regard $\hat{p}$ as a certain bound [^2] Let $p_{+}=\hat{p}$ if $\hat{p}>p$, else $p_{-}=\hat{p}$. of CI and substitute these variables into Eq. (1), we can solve the deviation score $z\in[0,+\infty)$ using numerical algorithms, such as the bisection method. A larger $z$ indicates a greater bias between model predictions and current observations, but a higher confidence level as well. We can directly use $p$ as the calibrated score, i.e., change $z$ to 00, to entirely eliminate biases. However, this will minimize the confidence level and result in poor generalization under data sparsity. Therefore, we intend to make the calibrated score between original $p$ and predicted $\hat{p}$, thus corresponding to a relatively high confidence. To this end, we propose a non-linear transformation function $g(z)$ to reduce the deviation. Intuitively, $g(z)$ should satisfy the following requirements: (1) passing through the origin: $g(0)=0$; (2) monotonic: $\forall z_{1}\leq z_{2},g(z_{1})\leq g(z_{2})$; (3) bounded: $\exists C\in\mathbb{R}^{+},g(+\infty)\leq C$. Thus, we devise a variant of the sigmoid function as follows:
@@
where $\lambda$ is a hyperparameter used to provide an upper bound and control the calibration intensity manually. A larger calibration intensity can result in a smaller $z$, meaning more belief in observations and less belief in original predictions. Obviously, we have $g(+\infty)=\lambda$ so that a large deviation is calibrated to a small one (e.g., the deviation corresponding to 95% confidence level if $\lambda=1.96$), and minor deviations are reduced almost linearly. By replacing the original $z$ in Eq. (1) with the transformed value $g(z)$, we can obtain a new CI, then one of its bounds is the desired confidence-aware calibrated score $\hat{p}^{\prime}$. In this way, given a test sample located in the same subset, we compute a field-wise calibrated score $p_{\text{calib}}$ by scaling its original predicted probability $p_{\text{pred}}$, i.e.,
@@
Overall, our confidence-aware method adjusts model-predicted probabilities by resetting the confidence level, which considers the observation and prediction together to avoid violent value changes and overfitting limited observations. Consequently, it keeps robust under severe data sparsity, especially circumstances with few observed samples and zero positive samples.
@@ 3.3. Multi-field Joint Calibration
Previous field-level calibration methods only calibrate and evaluate errors in a certain field, ignoring other fields. However, the field-level calibration error in different fields may vary greatly, which is ignored by single-field calibration. In addition, the biases in different fields, as reflections of the global bias from different perspectives, have a certain degree of commonality. Therefore, data sparsity faced by a certain field is expected to be alleviated through calibrations on other fields.
A naive way to extend calibration from a single field to multiple fields is to split sample subsets based on the combination of feature values of all the fields. Nevertheless, it will cause severer data sparsity since fewer samples satisfy multiple field value conditions. Accordingly, prior single-field calibration methods cannot be directly extended to multi-field calibration. In contrast, our method demonstrates a better extensibility among fields. The confidence-aware calibration in a single field in Section 3.2 is actually a scaling operation for a specific sample, thus allowing a generalization to multi-field joint calibration. Assume that there are $K$ target fields to calibrate. For each sample, we can obtain $K$ scaling multipliers $m_{1},m_{2},\dots,m_{K}$, where $m_{i}={\hat{p_{i}}^{\prime}}/{\hat{p_{i}}}$ for the $i$-th field. We adaptively fuse these multipliers via the weighted geometric mean, that is
@@
where $w_{1},w_{2},\dots,w_{K}$ is weights satisfying $\sum_{i=1}^{K}{w_{i}}=1$. The fusion weights can be precisely obtained by grid search. The proposed fusion method allows us to conduct joint calibration based on the marginal distribution of each field independently instead of dealing with the joint distribution, thus avoiding further sparsity of samples.
Figure 4. Online deployment diagram of ConfCalib.
@@ 3.4. Online Deployment
Fig. 4 shows the online deployment diagram of ConfCalib in our advertising platform. On the offline side, the prediction model’s training and ConfCalib’s fitting are paralleled and completely independent. At fixed intervals, the calibration module pulls the latest online samples from the log database and refits calibration mappings. This interval is usually set relatively shorter (e.g., 30 minutes) than the prediction model update period to better capture the dynamic shift of online data. On the online side, prediction and calibration are conducted in sequence. First, the prediction model makes predictions for candidate objects, then the calibration model calibrates the predicted probabilities and outputs the calibrated probabilities for subsequent use.
@@ 4.1.1. Settings
For all the datasets and methods, the prediction model is a fully connected network with four layers. Hyperparameters of baselines are set according to the original papers or carefully tuned. Results of neural methods are the average results of five random experiments with means and standard deviations reported. For our ConfCalib, we set the number of bins as {50, 10} and $\lambda$ in Eq. (2) as {1.0, 1.6} for Avazu and our industrial dataset.
Table 1. Data statistics of the public and industrial datasets.
@@
4.1.2. Datasets
We conduct experiments on our industrial dataset and a public dataset Avazu  [^3] https://kaggle.com/competitions/avazu-ctr-prediction. Data statistics are shown in Table 1. Avazu is a popular dataset on mobile ad CTR prediction with 40.4 million samples in 10 days ordered by the impression time. We split the dataset by day in a ratio of 7:1:2 for training, validation, and test. For our multi-field joint calibration method, the target fields are set to ”site_id” (3480) [^4] Represent the number of possible values of the corresponding field., ”app_id” (4864) and ”app_domain” (305). For other methods considering a single field, the target field is set to ”site_id”. The industrial dataset is collected from the click logs on our ad platform in 10 consecutive days, and the prediction target is conversion, with a split ratio of 7:1:2 by day. For our method, the target fields are ”advertiser_id” (529), ”app_size” (82), and ”app_id” (56). For other methods, the target field is ”advertiser_id”. The ratios of validation sets are relatively small while splitting, which is more in line with the short time window of online data collection.
4.1.3. Baselines
We choose several representative methods of the three categories as baselines for comparison. The scaling-based methods include naive scaling (multiply all predicted scores by a value $k$ to make their mean equal to the positive sample ratio), Platt Scaling (PlattScaling) (Platt et al., 1999), Gaussian Calibration (GaussCalib) (Kweon et al., 2022) and Gamma Calibration (GammaCalib) (Kweon et al., 2022). The binning-based methods include Histogram Binning (HistoBin) (Zadrozny and Elkan, 2001) and Isotonic Regression (IsoReg) (Zadrozny and Elkan, 2002). The hybrid methods include Smoothed Isotonic Regression (SIR) (Deng et al., 2021), Neural Calibration (NeuCalib) (Pan et al., 2020) and AdaCalib (Wei et al., 2022), with the latter two methods based on neural networks. The number of bins for NeuCalib and AdaCalib is set according to the initial settings when proposed.
4.1.4. Evaluation Metrics
Referring to prior methods, we continue to use field-level relative calibration error (Field-RCE or F-RCE) (Pan et al., 2020) to evaluate the calibration error on a certain field, which formula is
[…]
where $\mathcal{D}$ is the dataset, $\mathcal{Z}$ is the possible value set of the target field and $\mathcal{D}_{z}$ is the subset of samples with the target field value $z$ satisfying $\sum_{z\in\mathcal{Z}}{|\mathcal{D}_{z}|}=|\mathcal{D}|$. Intuitively, a lower Field-RCE indicates lower biases along the target field. Furthermore, we use the mean Field-RCE on multiple fields (Multi-Field-RCE or MF-RCE) to measure the multi-field calibration performance.
where $R$ is the number of shuffle operations, $T$ is the number of bins, and $\mathcal{D}_{r,t}$ is the subset of samples in the $t$-th bin after the $r$-th shuffle. MVCE allows us to inspect the calibration error from a more global perspective beyond the level of feature fields, thereby highlighting the effectiveness of multi-field calibration.
Additionally, we also report AUC and LogLoss to evaluate the ranking performance after calibration since our method and primary baselines are not globally isotonic.